What is the standard deviation of each stock?

Suppose that the index model for stocks A and B is estimated from excess returns with the following results:

RA = 2.2% + 0.80RM + eA

RB = –2.2% + 1.2RM + eB

σM = 24%; R-squareA = 0.16; R-squareB = 0.12

What is the standard deviation of each stock? (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Standard Deviation
Stock A %
Stock B %
Answer #1

Solution :-

RA = 2.2% + 0.80 RM + e A

RB = –2.2% + 1.2 RM + e B

σM = 24%; R-square A = 0.16; R-square B = 0.12

Now from the equation

Beta of Stock A = 0.80

And Beta of Stock B = 1.2

R-square A = Beta A * Beta B * σM * σM / (σA * σA) =

0.16 = 0.80 * 1.2 * 0.24 * 0.24 / (σA * σA)

(σA * σA) = 0.3456

σA = Standard Deviation of A = 58.79%

R-square B = Beta A * Beta B * σM * σM / (σB * σB) =

0.12 = 0.80 * 1.2 * 0.24 * 0.24 / (σB * σB)

(σB * σB) = 0.4608

σB = Standard Deviation of B = 67.88%

If there is any doubt please ask in comments